OpenStudy (anonymous):

Find the domain of the function f(x)= 3x+1/√x^2+x-2?

OpenStudy (anonymous):

\[(1,\infty)\]

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

if this is \[f(x)=\frac{3x+1}{\sqrt{x^2+2x-2}}\]

OpenStudy (anonymous):

(-infinity,-2) sry

OpenStudy (anonymous):

you have to make sure that \[x^2+x-2>0\]\[(x+2)(x-1)>0\] zeros are at -2 and 1, and since this is a parabola facing up it is negative between the zeros and positive outside. your answer is \[(-\infty,-2)\cup (1,\infty)\]

OpenStudy (anonymous):

yes what satellite said